Find the effective annual interest rate for each case. (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

  

APR Compounding Period Effective Annual Rate   

12% 1 month ____%
8% 3 months ____%
10% 6 months ____%
Answer #1

EAR = [1 + (APR / m)]m – 1

EAR = [1 + (0.12 / 12)]12 – 1 = 0.1268 or 12.68%

EAR = [1 + (0.08 / 4)]4 – 1 = 0.0824 or 8.24%

EAR = [1 + (0.10 / 2)]2 – 1 = 0.1025 or 10.25%